Mr. Toriop owns 5,000 shares of stock in Yummy Corporation. The company has announced that it will pay a dividend of $5 per share in one year and a liquidating dividend of $50 per share in two years. The required return on Yummy stock is 12%

c. Mr. Toriop wishes to have equal amounts of dividend income for the next two years. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends. (Hint: Dividends will be in the form of an annuity.)

d. Suppose Mr. Toriop is thinking about buying a house for $220,000 in one year. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends.

e. Suppose Mr. Toriop is thinking about postponing the house purchase for two years, by which time the price of the house will have increased by $46,800. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends.
Answer #1

Answer:

c. FV = 0, PV = 44.32, rate = 12%, n =2

use PMT function in Excel

amount of dividends = 26.41

d. total value per share in one year = 220,000/5000 = 44

value today = 44/1.12

FV = 0, PV = 39.2857, rate = 12%, n =2

use PMT function

home made div = 23.25

e.

FV = 0, PV = (220,000+46800)/(5000*1.12^3), rate = 12%, n =2

use PMT function

home made div = 22.47